### v1

# Rheological Analysis and Evaluation of Measurement Techniques for the Curing Polymethylmethacrylate Bone Cement in Vertebroplasty

Abstract
Vertebroplasty is a minimally invasive surgical procedure used to treat vertebral fractures, which conventionally involves injection of a polymethylmethacrylate (PMMA) bone cement into the fractured vertebra. A common risk associated with vertebroplasty is that of cement leaking out of the vertebra during the injection. This could happen since the flow behaviour of the bone cement is not conventional, but it varies with many factors like curing time, shear rate, and temperature. Therefore, experiments to quantify the cement’s flow properties are necessary to be able to understand and properly handle the bone cement. In this study, we aimed to characterize the behaviour of the PMMA bone cement in its curing stages, in order to obtain parameters that govern the flow behaviour during injection flow. The measurements were done using rotational and oscillatory rheometry as well as using a custom-made injetor setup replicating a typical vertebroplasty setting. Our results showed that the complex viscoelastic behaviour of the bone cement, which evolves over a two-phase curing process, can be significantly affected by deformations and temperature. We found that the results from rotational tests were affected by wall slip and “ridge”-like formations in the test sample, which we confirmed visually by microscopical assessment. We also found that the Cox-Merz rule to be conditionally valid, which impacts the use of oscillatory tests to obtain shear-thinning characteristics for the bone cement.
keywords:
[…]
1 Introduction
Vertebral fractures are a major source of suffering and disability, carrying risk of back pain, immobility, and potentially severe post-complications that impact the quality of life 1, 2. Up to 28% of patients with osteoporotic fracture-related issues resulted in death by 2008 3. Unlike many other musculoskeletal complications where bone grafting or alternative methods can be used 4, the most commonly used procedure to treat vertebral fractures is vertebroplasty 5, 6, 7. In this procedure, a rapidly curing bone cement is injected into the porous cancellous part of the vertebra to serve as mechanical support 8, 9, 10 and allow early mobilization, especially in the elderly, where lack of mobility often causes a cascade of medical issues and often death. A successful vertebroplasty can provide instant pain relief 11 to the patient. On the other hand, improper injection could lead to cement leaking outside the vertebra, which could cause serious complications like pulmonary embolism and paralysis 12, 13. The bone cement is a material with complex mechanical behaviour dependent on many factors, which makes it hard to predict its behaviour during injection, especially for practitioners who rely on intuition, skill, and sensory feedback. Therefore, characterization of the curing bone cement using various measurement techniques is helpful to understand its behaviour during injection in various conditions. Mathematical constitutive models derived from the characterization data can quantify the various dependencies, which can be used in analytical or computational models to reliably predict the behaviour of the bone cement during injection 14, 15, 16, 17, 18, 19.
The bone cements typically consist of ceramic components and fast-curing, minimally-toxic polymers that rapidly polymerize into, e.g. poly(methyl methacrylate) (PMMA) 8, 20, 21, 22. The rapid polymerization (curing) of bone cement is designed to provide the required structural support for the damaged vertebra, whereas ceramic components like zirconium dioxide function as radio-opacifiers that allow for easy visualization during the vertebroplasty procedure 23. Hydroxyapatite, a calcium phosphate mineral naturally found in bone, is also sometimes added to improve biocompatibility and to promote bone growth around the cement. The curing typically relies on free radical polymerization, the initiators react to form free radicals which cause the methlyl methacrylate (MMA) monomer molecules to form PMMA polymer chains 24, 25. The polymerization is an exothermic process 24, meaning that it is further accelerated by the heat released and the energy introduced, including mechanical shear.
In general, it should be noted that PMMA bone cement is far from behaving like an ideal fluid, rather it has a dough-like consistency while it is curing. The behaviour of the bone cement can be categorized as viscoelastic, i.e. it is a combination of solid-like and fluid-like behaviour. Moreover, the behaviour is dependent on many factors, including time, flow conditions (e.g. shear rate), and temperature. Due to this multifactorial dependency, the complete characterization of the bone cement is a challenge. There are many previous studies measuring the rheological properties of the bone cements. Different studies use different brands of PMMA bone cement for their measurements. Moreover, the set of conditions used to carry out the measurements are different in each study. Not only the conditions, but also the methods used for the measurements are different in each study, e.g. while some studies report steady shear viscosity from rotational shear measurements, others report complex viscosity from oscillatory measurements. The steady state and complex viscosities can be compared only using the Cox-Merz rule 26, which says that the complex viscosity measured must be equal to the steady shear viscosity if the angular frequency is equal to the shear rate, respectively. However, this rule is not valid for all kinds of materials. Nevertheless, its validity for the PMMA bone cement is often assumed rather than being explicitly investigated. The differences in some of these studies are shown in Table 1. Due to these reasons, it is difficult to compare or compile the data from these studies to understand the complete rheological behaviour of a bone cement over the range of shear rates and temperatures expected during a vertebroplasty.
Hence, the aim of this work was to understand the flow behaviour of the bone cement during injection in its curing phase in conditions relevant to vertebroplasty, taking Vertecem V+ as an example. The main objectives of this work were as follows:
1.
To quantify various aspects affecting injection flow behaviour, viz. viscosity, yield stress, loss factor, flow point, etc. from measurements
2.
To fit the experimental results to mathematical models that describe the rheological behaviour, like the power law model, and obtain the model parameters that can be used for prediction and simulation of injection flow behaviour
3.
To compare the different methodologies used for rheological measurement, viz. injection tests as well as rotational and oscillatory measurements on the rheometer, namely their advantages and limitations
4.
To check the validity of the Cox-Merz rule for the PMMA bone cement
| Authors | Cement | Methods | Measurements and conditions |
| Boger et al. 27 | Unnamed | Oscillatory | Time evolution, 1 Hz, various |
| Deusser et al. 28 | Vertecem | Oscillatory | Time evolution, 1 Hz, 23 °C and 37 °C |
| Dunne and Orr 29 | Unnamed | Capillary | Time evolution |
| Farrar and Rose 30 | Various | Oscillatory | Time evolution, 0.05%, 5 Hz, various |
| Kolmeder and Lion 31 | Osteopal | Rotational CP | Time evolution, various |
| Capillary | Time evolution, various | ||
| Kolmeder et al. 32 | Unnamed | Oscillatory PP | Time evolution, 1 Hz, 0.1%, various |
| Krause et al. 33 | Various | Rotational CP | Time evolution, 0.04 and 1 s |
| Capillary | Shear rate dependence, 23 °C | ||
| Lepoutre et al. 17 111These studies assume the validity of the Cox-Merz rule. | Osteopal | Oscillatory PP | Time evolution, 50%, various and |
| Injector | Shear rate dependence, various | ||
| Lewis and Carroll 34 | Orthoset | Oscillatory CuP | Time evolution, 1 Hz, 18 °C |
| Lian et al. 18 111These studies assume the validity of the Cox-Merz rule. | Simplex | Oscillatory PP | Frequency sweep, 0.2% |
| Nicholas et al. 35 | Various | Oscillatory PP | Time evolution, 5 Hz |
2 Experimental
2.1 Bone cement
The bone cement used for this study was purchased as non-sterile bulk material from OSARTIS GmbH (Germany), which is equivalent of the bone cement commercially sold by the name VertecemTM V+ by DePuy Synthes. The bone cement is supplied as a powder consisting of 40% zirconium dioxide, 15% hydroxyapatite, 44.6% poly(methyl methacrylate) (PMMA), and 0.4% benzoyl peroxide; and a liquid consisting of 99.35% methyl methacrylate (MMA) stabilized with 60 ppm hydroquinone, and 0.65% N,N-dimethyl-p-toluidine (DMPT). Typical application of the bone cement uses the provided cement mixing kit (DePuy Synthes), where the initiator and powder are mixed with MMA monomers in 26 grams of powder to 10 mL of liquid ratio, upon which immediately the user is required to push and pull the mixing chamber kit handle from endpoint to endpoint for 20 seconds, at a steady pace of 1–2 strokes per second. The resultant mixture has a dough-like consistency which can be filled into syringes and injected. The polymerization occurs through the free radical polymerization method, wherein the activator DMPT reacts with the initiator benzoyl peroxide to form free radicals. These free radicals react with monomer MMA molecules that cause them to form polymer chains of PMMA. PMMA is the main component of bone cement and is responsible for its mechanical properties. The entire polymerization process takes about 20 minutes at room temperature.
2.2 Rheometer measurements
The rheological measurements were performed on MCR302 rheometer from Anton Paar using a parallel plate (PP) setup. Disposable top and bottom plates were used for the measurements. The diameter of the top plate was 25 mm. The bone cement contains particles of size 5.2 ± 1.5 microns 36, hence the gap between the two plates was kept at 1.5 mm to ensure sufficiently larger gap size compared to the particle size 30 and avoid any particle-scale effects. The temperature during the measurements was controlled using the Peltier module of the rheometer, so that the measurements are not affected by the heat released during curing. The temperature was fixed at 23 °C in all tests, unless otherwise specified. The rheometer and the associated instrumentation was switched on for at least one hour to equilibrate the temperature and all other conditions before carrying out measurements.
The bone cement prepared using the typically provided mixing kit would produce a large offset of unnecessary material for measurements. Hence, we established a mixing process for our laboratory conditions that produced only sufficient amount of bone cement for each measurement. To ensure that our laboratory mixing process replicates the bone cement behaviour from the original mixing process, we established a benchmark test as quality control. In the test, the bone cement sample was subjected to oscillations at a constant torque of 3 mNm and 1 Hz frequency at 23 °C, with a measuring point every 5 seconds. After multiple trials, the successful and reproducible method consisted of the following steps:
1.
2.6 grams of the PMMA powder were weighed in a 10 mL beaker.
2.
Separately, 10 mL of the MMA monomer liquid was prepared in a batch glass vial.
3.
At time = 0 s, 1.0 mL MMA monomer liquid was dropped into the 10 mL beaker with the powder using a positive displacement pipette, and a stopwatch for measuring the time was started.
4.
The powder and the liquid were stirred using a non-reacting polyetheretherketone (PEEK) stirring road for 20 seconds and counting about 20 rotations.
5.
After 20 s, part of sample was gently dropped onto rheometer bottom plate and top plate was gently lowered down to avoid any disturbance.
6.
Silicone oil was then spread around the sample and the humidity control hood was used to avoid sample evaporation. The test was then started and the time on the stopwatch was recorded.
The test on the rheometer commenced once the bone cement was prepared and placed between the parallel plates. The time since start of mixing to start of test for all tests varied in the range of 180 30 seconds. For the characterization of the bone cement, various tests were done using the rheometer, namely
Test Rh1a: Oscillatory test at 0.2% strain amplitude and 1 Hz frequency for 30 minutes: this test was done to observe the evolution of the bone cement properties in time at low strain value.
Test Rh1b: Oscillatory test at 20% strain amplitude and 1 Hz frequency for 30 minutes: this test was done to observe the evolution of the bone cement properties in time under continuous larger deformation.
Test Rh2a: Oscillatory test with 0.1% strain amplitude, 0.1 Hz (5 mins) – 20% strain amplitude, 1 Hz (3 mins) – 0.1% strain amplitude, 0.1 Hz (22 mins). This procedure was done to replicate a typical injection, in which the bone cement is at rest at first after mixing, then it is deformed as it is filled and injected through the syringe and cannula, and finally is at rest once it is injected. The aim here was to investigate how the bone cement breaks under deformation and recovers thereafter.
Test Rh2b: Oscillatory test with 0.1% strain amplitude, 0.1 Hz, 23 °C (5 mins) – 20% strain amplitude, 1 Hz, 23 °C (3 mins) – 0.1% strain amplitude, 0.1 Hz, 37 °C (22 mins). This procedure was done to replicate a typical injection into a human vertebra, hence the change in temperature in the last step from room temperature to physiological body temperature.
Test Rh3: Rotational shear stress applied in steps of 0 Pa (1 min) - 100 Pa (1 min) - 0 Pa (1 min) - 500 Pa (1 min) - 0 Pa (1 min) - 2000 Pa (1 min) - 0 Pa (1 min) to evaluate creep behaviour.
Test Rh4: Rotational shear strain applied in steps of 0% (2 min) - 1% (2 min) - 0% (2 min) and 0% (2 min) - 100% (2 min) - 0% (2 min) to evaluate stress relaxation behaviour.
Test Rh5: Amplitude sweep from 0.01% to 1000% strain amplitude, 1 Hz frequency to obtain the yield stress and the crossover point of storage and loss moduli.
Test Rh6: Frequency sweeps from 0.1 Hz to 100 Hz at various strain amplitudes. The frequency sweep yields the dependence of the complex shear modulus and the complex viscosity on oscillation frequency. The strain amplitudes were chosen a posteriori as per the different regions obtained from the amplitude sweep test.
Test Rh7a: Rotational shear rate sweep, various tests were done to obtain flow curves with the shear rate ranging over 2 or 3 orders of magnitude in each test. In this way, steady-state viscosity values for shear rates from 10 s to 10 s were obtained.
Test Rh7b: Rotational shear rate sweep from 0.1 to 100 s with 100-grit sandpaper stuck to the top and bottom plates using double-sided tape. The test was done to check if there was a difference in the result with and without sandpaper, which would provide evidence for the occurrence of wall slip during the test.
Test Rh8: Rotational tests with constant shear rate for 20 minutes. These tests are done at shear rates 10, 10, 10, 10, 1, and 100 s. These tests were done to evaluate the time-evolution of viscosity at various flow rates.
2.3 Optical microscopy (tests Rh9 and Rh10)
Two additional tests were done on the rheometer to make bone cement samples for investigation with optical microscopy. In the first test (test Rh9), bone cement was prepared and simply left on the rheometer between the plates while maintaining 1.5 mm gap for 45 minutes without any action. In the second test (test Rh10), bone cement was prepared and subjected to 15 minutes of rotational 100 s shear rate, like in test Rh7b, and then 30 minutes of no action on the rheometer. The two cured bone cement discs formed as a result of these tests were investigated using an Axiotech microscope, which is a material microscope for non-transparent samples, equipped with a 2.5x objective. The images were captured using a digital camera Axiocam 105, and processed using software Axiovision 4.9.1. The microscope, camera, and processing software were provided by Carl Zeiss AG, Germany. Reflected light bright field technique was used for the illumination.
2.4 Tests on the custom-made injector (Tests Inj1 – Inj5)
The setup for the injection experiments is shown in Figure 0(a). The injector consisted of a carriage which was driven by a stepper motor using a ball screw with 5 mm feed per revolution. The stepper motor had a resolution of 1.8° corresponding to 200 steps per revolution, amounting to a calculated carriage resolution of 0.025 mm. The syringe was placed horizontally through a hole at the other end of the device as shown in the Figure 0(a) and held in place with a bolt. The cannula was then attached at the end of the syringe. The carriage pushed the plunger at the desired rate which could be programmed using a computer. At the point of pushing the plunger, a 200 N load cell was mounted on the carriage to measure the forces applied to the plunger during injection.
To prepare the bone cement, a 10 mL syringe was taken and sealed at its open end using a removable plug. The plunger of the syringe was removed and 3 mL of the monomer liquid was poured in it using a pipette. 7.8 grams of the PMMA powder were added directly into the syringe using a funnel and then mixed in the syringe using a PEEK stirring rod for 20 seconds (20 rotations). After mixing the components, the plunger was put back in the syringe and the plug at the other end was removed. The mixed bone cement was then first transferred to a 2 mL syringe until it was completely full, which was then used to fill up an 8-gauge cannula. The cannula had a side-opening and open front end, of which the side-opening was sealed using a strong adhesive tape so the bone cement could only come out of the front end when injected. The emptied volume of the 2 mL syringe was then again filled up using the remaining bone cement in the 10 mL syringe. At the beginning of the test, the carriage was set at the ‘unloaded’ position, which is far back to be able to mount the filled 2 mL syringe easily. Once the syringe and the cannula were mounted, the carriage was set to ‘loaded’ position, which was just behind the fully pulled back plunger. The ‘unloaded’ and ‘loaded’ positions were programmed in the device so it automatically arrived at them when commanded. At this point, the device was given the command to inject with a specified speed according to the required flow rate. The time from the start of mixing to the start of injection was measured using a stopwatch. The tests, named Inj1 to Inj5, were carried out with flow rates of 0.025, 0.05, 0.1, 0.2, and 0.4 mL/s respectively. During the injection, the load cell measured the force required for injection every 0.05 seconds. The time from mixing to start of injection in all tests was between 230 and 245 seconds.
2.5 Analytical calculations
The measurements obtained from the injection tests were then used to obtain the rheological parameters of the bone cement according to the power law constitutive equation
where is the viscosity, is the shear rate, and consistency index and flow index are the material parameters. Assuming such a fluid flows through a tube of radius and length with flow rate , the shear stress at the wall relates to the pressure difference as
Furthermore, the shear rate in the tube can also be obtained using the equation
Detailed derivations of these equations can be found in the works of 37, 38.
To apply these equations to our problem, the geometry of the apparatus was simplified as shown in Figure 0(b). The syringe and the cannula were assumed to be cylinders with the given dimensions. The coupling connecting the two parts had the same inner diameter as the cannula, hence its length was included in the cannula in the simplified geometry. The nozzle of the syringe was short in length and was therefore ignored. The change of cross-section from syringe to cannula would add to the pressure loss in reality, however this was ignored in our study. We then used the Equation 3 to obtain the pressure loss in our system. The pressure loss is then the sum of losses in the syringe and the cannula , given as
where mm, mm were the length and radius of the syringe respectively. Similarly, mm and mm were the length and radius of the cannula respectively. Given force equilibrium in the system, the pressure loss obtained from Equation 5 is the pressure that would be required to be created by the injection force applied externally on the syringe plunger. Hence, the required injection force is then given as
3 Results and Discussion
3.1 Oscillations at constant strain and frequency (Tests Rh1a and Rh1b)
At 0.2% strain amplitude and 23 °C, the complex modulus of the bone cement increased with time due to the ongoing polymerization. The polymerization occurred with a slow first phase, when the modulus increased gradually at a rate of about 29 kPa/min (Figure 1(a)). This is the working phase when the bone cement gradually hardens with time but maintains an injectable paste-like consistency. The slow first phase was followed by a rapid curing phase after about 20 minutes when the bone cement polymerized rapidly, and became solid in the next 5-10 minutes. The complex modulus after solidification was above 25 MPa. Interestingly, the loss factor always remained below 1, implying that the bone cement shows elastic prevalence, i.e. behaving more like an elastic solid than a liquid, even at the early stages of curing. The loss factor was between 0.5 and 0.7 until 25 minutes and rapidly decreased thereafter (Figure 1(a)).
When the same test was carried out at higher 20% strain amplitude, the modulus was about two orders of magnitude lower (Figure 1(b)). Zooming in on the graph showed that initially, the modulus decreased with time. The decrease was an exponential decay at first and a continued decrease afterwards. The exponential decay is typical of viscoelastic stress relaxation behaviour, while the continued decrease afterwards could be probably because of fatigue resulting from repeated oscillations. After about 8 minutes, the modulus started to increase again, which was probably due to new polymer chains forming faster than the rate of breaking. The increase was, however, much slower compared to that observed in the test with 0.2% strain. Interestingly, the loss factor here was greater than 1, implying that the bone cement behaves more like a fluid when it is being significantly deformed. The loss factor did not go lower than 1 until after 25 minutes, when the rapid curing phase had already begun.
The properties of the bone cement continuously evolve because of the curing reaction. Hence, it is important to evaluate the so-called mutation time, which is defined as the time required for change in a given property by a factor equal to the Euler’s number . Taking the complex modulus as the property of interest, the bone cement has a mutation time of about 180 seconds and 700 seconds at 0.2% and 20% strain amplitude respectively. For reliable measurements, it is important that the measurement times are much less than mutation times to avoid the effect of the evolving properties in measurements. This is ensured to be the case for most of our measurements. As an exception, the measurement times are not negligible in the initial few points in the frequency sweep but they are still much smaller than mutation times.
These results show that although the bone cement behaviour is predominantly solid when it is at rest or at very small deformations, while high deformations cause the bone cement to behave more like a fluid. This indicates that the bone cement is suitable for injections, despite this not being apparent from the test Rh1a. Furthermore, the curing process is inhibited by the deformations, as is evident from the much lower magnitude as well as a delayed and lower rate of increase in the complex modulus during the entire duration of the test.
3.2 Replication of injection (Tests Rh2a and Rh2b)
Figure 1(c) shows the results of the test Rh2a. During the initial stage of the test when the strain amplitude and frequency were low, the complex modulus increased gradually due to polymerization, slightly faster than at initial time of Rh1a (Figure 1(a)) test due to the lower strain used. In the next stage of the test, while it was subjected to higher strain and frequency, the modulus immediately dropped to 1% of the original magnitude and the loss factor jumped to a value higher than 1. This again showed the bone cement’s transition to more fluid-like behaviour as soon as it is deformed during polymerization. This was followed by an exponential decay of the modulus, as was also observed in the previous test Rh1b (Figure 1(b)). When the bone cement is brought to (almost) rest again in the final stage of the test (i.e. at about 12 minutes time), there is an immediate but small recovery of about 8% in the modulus. The polymerization is resumed and the rapid curing phase occurs at about similar time of 25 minutes, after which the bone cement solidifies. We note however that properties of bone cement modulated in this way (simulating more injection-like scenario), are different to those measured at low deformations in test Rh1a (Figure 1(a)). For example at the end of the test, the complex modulus is 2.5-fold lower and loss modulus is about 2-fold higher due to structural disturbances during curing. Carrying out the same test with the bone cement kept at 37 °C in the last stage of the test (Rh2b), there is hardly any recovery, however the rapid curing phase of the bone cement is triggered immediately. The bone cement is a fully cross-linked polymer solid at 22 minutes, as evident from the plateaued complex modulus and the loss factor value of 0.05.
The results show that the curing time is shortened when the bone cement is exposed to higher temperature. Once inside the vertebra of the patient, the bone cement starts curing rapidly. Vertebroplasty is often done in steps of smaller injections, in which case it is important for practitioners to be aware that the bone cement cures rapidly when exposed to the physiological body temperature. If the subsequent injections are not done in a timely manner, the viscosity of the cement inside the vertebra could get much higher than the one being injected. This viscosity difference could result in disadvantages like requiring higher injection force, unintuitive injection force development over the course of the injection, and unstable flow patterns inside the vertebra, as was highlighted in a previous work 19.
3.3 Step stress (Test Rh3) and step strain (Test Rh4)
Figure 2(a) shows the behaviour of the bone cement when subjected to step stress. The resultant strain was observed to be exponentially decaying, which is typical of viscoelastic behaviour, both when stress was applied and then removed. Upon removal of the stress, the strain did not converge back to zero, which indicated that there was permanent deformation in the sample. At very high stress, i.e. at 2000 Pa, the bone cement showed almost purely viscous behaviour as it deformed continuously, and there was no recovery after removal of the stress.
Figure 2(b) shows the stress response of the bone cement to step strain. Two tests were carried out with 1% and 100% step strain respectively. The resultant stress was maximum at the start of the applied step strain, and then decreased exponentially with time to a non-zero stress value, typical of viscoelastic stress relaxation behaviour. As the applied strain was removed, returning to zero strain state required stress in the opposite direction, which again showed similar relaxation behaviour. The decay curve could be fit to the relaxation equation
where is the time since the start of step strain, is the stress at time , is the stress at the start of relaxation, is the final stress the material relaxes to, and is the relaxation time. The relaxation times obtained from curve fitting are shown in Figure 2(b). Compared to the case of 1% strain, the relaxation times in the case of 100% step strain are lower, indicating that the bone cement relaxes faster owing to lesser resistance to deformation. This reinforces our previous observation that the bone cement has more fluid-like behaviour while it is being deformed.
3.4 Amplitude sweep (Test Rh5)
Figures 2(c) and 2(d) show the range and mean of the complex modulus, loss factor, and complex viscosity for three amplitude sweep tests on the bone cement. The modulus grew initially, likely because the strain was too small to inhibit the polymerization. At 0.02 - 0.04% strain, the modulus stopped growing and started lowering afterwards. The stress at which the complex modulus and the complex viscosity start to lower is the yield stress, which was observed to be around 70-110 Pa for this bone cement. This is the stress at which the bone cement starts to flow. Hence, this parameter is important to understand the injectability of the bone cement. The given yield stress range corresponds to injection force of about 0.7 N according to the equation obtained from Equation 2, which is relatively small and hence corroborates the use of bone cement in clinical injections. However, it is important to note that this value holds only for 295 20 seconds time since mixing, and would increase the more it is allowed to rest. In the 1-10% strain amplitude range, the loss factor went above 1. This is called the crossover point, or the flow point, and corresponded to about 500-1000 Pa stress ( 3-7 N of injection force). Finally, there was a sudden drop in the modulus near 200% strain or 2500-4000 Pa stress mark, which was likely because of insufficient friction between the rheometer plates and the bone cement, causing the plates to slip. The evidence for this phenomenon, also called wall slip, is provided in later sections.
3.5 Frequency sweep (Test Rh6)
Figures 2(e) and 2(f) show the range and mean of the complex modulus, loss factor, and the complex viscosity over three frequency sweep tests on the bone cement. We observed that the complex modulus kept increasing throughout the duration of the test. On the other hand, the complex viscosity initially increased slightly, then kept dropping as the angular frequency increased beyond 1.3 rad/s. This angular frequency can be considered the inception point of shear-thinning behaviour. The shear-thinning of the viscosity according to power law, i.e. Equation 1, started at about 2.6 rad/s. Fitting the curve to the power law equation yields the power law flow index . Value of between 0 and 1 indicates shear-thinning, which is expected for the PMMA bone cement. Measurements at higher angular frequency could not be executed due to the rheometer limitation of maximum frequency of 100 Hz.
3.6 Shear rate dependence (Tests Rh7a, Rh7b, and Rh8)
Figure 3(a) shows the results from the various steady shear rate sweep tests performed in rotational mode, also known as flow curves. The viscosity values were not affected by whether the test was done over a shorter shear rate ranges or all the way from 10 to 10 s. This implied that the duration of the test did not affect the results, at least within the first 10 minutes of mixing. The influence of curing was seen at very low shear rates in the order of 10 s, where the viscosity increased. The shear-thinning is steady until about 1 s. Upon curve-fitting, the flow index was obtained in the range of 0.20-0.30, and the consistency index in the range of 700-1100 Pa s.
As the shear rate went above 1 s, the viscosity drop was noticeably steeper, until it started plateauing at shear rates greater than 20 s. The values of were less than zero in this range, implying that the shear stress reduced as the shear rate was increased, as seen in Figure 3(b). This could be a result of structural breakdown of the material or a measurement artefact due to the suspected wall slip effect. Performing the same tests using sandpaper attached to both the top and bottom plates (see test Rh7b in ‘Experimental’ section), the results showed a slight improvement in the maximum shear stress achieved (Figures 3(a) and 3(b)). This provided evidence that the drop in shear stress indeed occurred due to wall slip. This was confirmed from visual inspection through microscopy, as detailed in a later section.
Figure 3(c) shows results of tests Rh8, i.e. time evolution at various constant shear rates. At the lowest shear rate 10 s, the bone cement showed a linear increase in viscosity with time of about 480 kPa s/min. The tests with higher shear rate showed lower magnitudes and slower increase with time. At 1 s shear rate, the viscosity kept decreasing for almost the entire duration of the test. At 100 s shear rate, the viscosity dropped almost entirely in the early stages (first 30 seconds) of the test . Viscosity values taken at various points in time for each of these tests were used to calculate the flow index with time. For this, the viscosity for shear rate 10 s was excluded since previous tests showed it was too low to induce shear-thinning. The results showed that the flow index was 0.33 in the beginning and 0.25 if the initial value from 100 s is taken into calculation. These values are in the same range as observed in tests Rh7a for shear rate less than 1 s. However, the value of kept dropping with time and became negative soon after the initial stages, as seen at 5 minutes mark in Figure 3(c). In fact, this was true even when the viscosity values from shear rates 1 and 100 s were excluded from the calculation for . This implied that the viscosity values only at the beginning of the test gave plausible shear-thinning characteristics. The same could be said for the consistency index, which when interpreted as the viscosity at unit shear rate, gave Pa s and kept decreasing with time.
3.7 Validity of the Cox-Merz rule for PMMA bone cement
According to the Cox-Merz rule, for the same magnitude of shear rate and angular frequency, the respectively measured steady shear viscosity and complex viscosity must also be equal 26. To extend this rule to materials with yield stress, Doraiswamy et al. 39 suggested multiplying the angular frequency to the strain amplitude for the Cox-Merz rule to be true. To validate this for the Vertecem V+ bone cement, the results from shear rate sweep and three repetitions of frequency sweep measurements at various strain amplitudes are superimposed in Figure 3(a). Note that, as we learned from the amplitude sweep measurements shown in Figure 2(c), the 0.01% strain amplitude is within the linear viscoelastic region, 0.03% is near the transition to non-linear region, 0.2% is in the non-linear region. The frequency sweep curves for 0.01% did not coincide with those of steady shear viscosity, but they were closer for 0.03%. The graphs coincided at 0.2%, which is in the non-linear viscoelastic range. As a check, a single frequency sweep was carried out at 200% strain amplitude, which is beyond the crossover point. The resulting curve seemed to extrapolate the curves from the tests with 0.2% strain, and coincided with the rotational shear rate graph until the aforementioned dip in viscosity due to wall slip. This is surprising since the frequency sweep is generally done at a strain amplitude that lies within the linear viscoelastic range. A possible explanation is provided by Shafiei-Sabet et al. 40, that the PMMA structures formed as a result of polymerization break under shear flow, but oscillatory tests at low strains do not affect these structures. This means that the Cox-Merz rule is not applicable to the Vertecem bone cement if the strain amplitude is low, i.e. within the linear viscoelastic range, but it is only valid for higher strain amplitudes in the non-linear range. This reveals the conditional nature of the validity of the Cox-Merz rule for the Vertecem bone cement.
3.8 Optical microscopy and visual inspection
Examples of cured bone cement samples after tests on the rheometer are shown in Figure 6. These samples were inspected microscopically to gain better insight into the results. Figure 6a shows an unsheared sample simply left at rest between the rheometer plates until it was cured, from test Rh9. Figure 6b shows sample subjected to 100 s for 15 minutes before letting it cure, from test Rh10. We observed circular “ridge”-like formations formed in the sheared sample as a result of continuous rotational deformation. The images from optical microscopy, shown in Figure 6e and 6f, showed that the depth of the ridges went almost through the entire thickness of the sample. The same ridges were observed even in the sample sheared at 0.1 s for 20 minutes (referred to test Rh8, Figure 3(c)), as shown in Figure 6c. A closer look, shown in Figure 6g, revealed that the sample looked normal in the center upto a radius of about 13.66 mm, and the ridges were formed only beyond this radius. In a parallel plate rheometer, the shear rate is not uniform across the sample, hence the specified shear rate is actually the average shear rate along the radius. Given that the shear rate at the periphery is 1.5 times the specified shear rate, and the shear rate increases linearly with the radius, we could calculate the critical shear rate as 0.082 s, beyond which the ridge formation was observed.
Interestingly, there were no visible ridges in the sample exposed to oscillatory deformation with 20% strain and 1 Hz for 30 minutes (refer test Rh1b, Figure 1(b)), as shown in Figure 6d. This implies that rotational shearing on the rheometer causes deformations in the bone cement that results in the formation of deep circular ridges, likely stemming from the combination of elastic-like instabilities and possibly Taylor-type vortex flows, as has been observed for other types of materials 41, 42, 43. This could happen even at low shear rates if the sample is sheared for sufficiently long time. The ridges would not only cause a loss of contact with the top plate, but also a lack of sufficient contact between neighbouring flow layers of the bone cement needed for viscous resistance, which would result in underestimation of the viscosity. This phenomenon limits the conditions in which the rheometer tests can be done on the bone cement and thus makes standardization of tests very difficult.
3.9 Tests on the custom-made injector (Tests Inj1 – Inj5)
The results of the injection experiments using the custom-made injector setup at various flow rates are presented in Figure 7, which shows the measured force against time normalized by the total duration of injection. As expected, the results showed that the higher the flow rate, the more injection force is required. For the two slowest flow rates, the force gradually increased because the duration of injection was long enough and the shear rates were small enough for the cement to cure while being injected. To obtain reliable results, the average force values for each flow rate over a 5 second time period closest to the 245 seconds mark from the time of mixing were used for the regression analysis, as shown in Figure 6(b). This was done to avoid any influence of curing time on the outcome of curve-fitting. The regression was carried out using Equations 5 and 6 obtained from analytical calculations, with power law parameters and as fitting parameters. The regression analysis gave a good fit with yielding values Pa s and . The value of is similar to those observed in test Rh7a for shear rate less than 1 s and in test Rh8 during the initial 30 seconds. Using Equation 4, we could calculate the maximum shear rate, which were 10 s and 165 s for the lowest and the highest flow rates respectively. For shear rates in this range, we observed decreasing viscosity with time in test Rh8 (Figure 3(c)). However, if this was truly the case, it would cause a decrease in the injection force with time, which was not observed in our results further evidence that the reducing viscosity seen in Figure 3(c) is due to the ridges formed in the sample as a result of the rotational deformation.
3.10 Comparison of the methods
| Test | Angular frequency * Strain amplitude/ | (Pa s) | |
| Shear rate / Flow rate | |||
| Rh6, 0.01% strain | 6.28 10 to 0.0628 rad/s | 331 | 0.33 |
| Rh6, 0.03% strain | 1.88 10 to 0.188 rad/s | 589 | 0.28 |
| Rh6, 0.2% strain | 0.00126 to 1.26 rad/s | 1060 | 0.27 |
| Rh7 | 0.001 to 1 s | 700-1100 | 0.20-0.30 |
| Rh8 | 0.001 to 1 s | 1083 | 0.33 |
| Inj1–Inj5 | 0.025 to 0.4 mL/s | 974 | 0.30 |
The values of the parameters of the power law equation and are useful for analyzing and simulating the injection flow behaviour of bone cement. Usually, these values are extracted from experimentally measured flow curves using rheometer tests like test Rh7a, or the oscillatory frequency sweep test Rh6 if the Cox-Merz rule is known to be valid. They can also be extracted from a capillary rheometer or an injector setup like in tests Inj1–5. The power law parameters extracted from these tests are summarized in Table 2.
Analytical calculations show that the shear rate range during an actual injection could go as high as 200 s. As previously discussed, the rotational rheometer is not suitable at such high shear rates because of the wall sip and ridge formation caused by rotational deformation. This could be avoided by reducing the gap size between the rheometer plates, but that cannot be done in this case since that the gap size must also be sufficiently large to ensure no particle effects. For the same reason, we could not use our cone-plate setup on the rheometer as it could only be used with a fixed gap of 49 microns, which is not large enough for our application. Hence, only measurements for shear rates below 1 s can be reliably used, as the power law parameters thus obtained are close to those obtained from the injection experiment. It must be noted, however, that even for low shear rates, the ridge formation could occur if the test duration was sufficiently long. It is important to exercise these These cautions when performing rotational rheometer tests on PMMA bone cements.
The flow curves can also be obtained from the oscillatory frequency sweep, but as seen from Figure 5, it is important to ensure that the strain amplitude used for the frequency sweep lies within the non-linear viscoelastic range. Only then is the Cox-Merz rule valid and a feasible comparison to steady shear flow (like during an injection) possible. This is a crucial finding since this is in contrast to the conventional norm of performing the frequency sweep with a strain amplitude within the linear viscoelastic region. Note that the viscosity drop seen as a result of wall slip in the steady shear viscosity from rotational measurements is not seen in the oscillatory frequency sweep results. This, along with previously detailed visual and microscopical inspection, confirmed that that problem of wall slip or ridge formation is absent in oscillatory measurements. This demonstrates the advantage of oscillatory tests over rotational tests, especially as the shear rates go higher.
The setup of a capillary rheometer or the custom-made injector setup used in this study is closest to an actual vertebroplasty setting. This kind of setup circumvents the drawbacks of the rotational or oscillatory rheometry, and yields mechanical parameters like effective or apparent viscosity and injection force that is most representative of the real-world setting. It is important to note that the flow model used and the analytical calculations play an important role in the accuracy of the extracted rheological parameters. The Herschel-Buckley model is arguably the more suited model since the bone cement exhibited yield stress. However, the yield stress is low compared to the shear stresses applied during the measurements in most cases, hence the power law model was used in this study for simplicity. There needs to be correction factors applied for flow entry, exit, change in tube diameter, etc., but they are beyond the scope of this work. Hence, the rheological parameters extracted from our setup are indicative numbers, and are not expected to have the accuracy or sophistication of industry-standard instruments.
4 Conclusion
From the various tests done, the curing PMMA bone cement was found to exhibit viscoelastic behaviour with a two-phase curing process: a slow initial phase followed by a rapid curing phase. The curing process is significantly affected by deformations and surrounding temperature. Exposure to physiological body temperature immediately triggered the rapid curing phase resulting in sudden increase in viscosity. This could negatively affect the vertebroplasty procedure if it is done in step-wise injections as a result of flow instabilities caused by injection of a low-viscosity fluid in a high-viscosity fluid environment. Properties like low yield stress and fluid-like behaviour under deformation, as opposed to predominantly elastic contribution at rest, make it suitable for injection.
While measuring the shear-thinning characteristics, rotational rheometry underestimated the viscosity above 0.082 shear rate. It was confirmed from microscopical visual inspection that the dip in viscosity was caused due to slippage between material and rheometer plates, and formation of circular ridges within the sample material. The same was not observed in a sample subjected to oscillatory rheometry and tests using our injector setup.
An important finding of the study is that Cox-Merz rule is valid only for oscillatory frequency sweep tests done at strain amplitude in the non-linear viscoelastic region, in contrast to the conventional norm of doing the test within the linear viscoelastic region. This conditionally validity of the Cox-Merz rule is important to take into consideration, especially when oscillatory measurements are used to obtain the flow curves.
Funding
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project Number 327154368 – SFB 1313. We also acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). J.K.W. acknowledges the European Union’s Horizon 2020 (H2020-MSCA-IF-2019) research and innovation programme under the Marie Sklodowska-Curie grant agreement 893099 — ImmunoBioInks.